# Agentic 3D Scene Generation with Spatially Contextualized VLMs

### v2

@@ 5 Experiments
Metrics. To assess the semantic alignment of the generated 3D scenes with the input prompts, we render images from the synthesized scenes and compute text-image similarity using CLIP [35] and BLIP [26], and image-image similarity using LPIPS (AlexNet) [65]. To further evaluate the aesthetic quality (i.e., realism and visual appeal) and functional plausibility (i.e., adherence to ergonomic principles), we collect human and GPT-4o ratings and report relative rankings across all methods. In Table 1, each method is ranked based on averaged ordinal scores across a benchmark set of scenes, with lower ranks indicating better performance. Human ratings are collected from a user study involving 16 participants.
Table 1: Quantitative comparison on 3D scene generation. Our method achieves the best performance across consistency (CLIP/BLIP), image fidelity (LPIPS), aesthetics (AQ), and functionality (FP).
Method CLIP ($\uparrow$) BLIP ($\uparrow$) LPIPS ($\downarrow$) AQ (4o/User) ($\uparrow$) FP (4o/User) ($\uparrow$)
@@ 5.1 Comparison
Image-conditioned generation. Figure 4 shows a comparison with ACDC [9], a recent method for real-to-sim scene construction. Our system more effectively reconstructs spatial layouts and scene compositions, such as the tilted sofa in a living room, while better preserving the stylistic integrity of iconic works like Van Gogh’s Bedroom in Arles. In Table 1, we report the best image-image similarity score, indicating higher visual fidelity to the input images. This advantage stems from our structured spatial context, which preserves geometric details and enables adaptive reconstruction.
Image set as input. Unlike prior methods, which are typically restricted to single-view inputs, our framework naturally accommodates unstructured and unposed image collections. As illustrated in Figure 6, our system consolidates geometric cues from diverse viewpoints into a coherent 3D layout. This ability stems from the VLM’s integration with our spatial context, which provides a flexible representation for resolving spatial correspondences across views.

### v3

@@ 5 Experiments
Metrics. (i) Geometric fidelity. We use Chamfer Distance (CD), which averages two terms: accuracy (the smallest Euclidean distance from reconstructed shape points to ground-truth points) and completeness (the smallest Euclidean distance from ground-truth points to reconstructed shape points). (ii) Instance overlap. We compute Intersection over Union (IoU) to measure instance-level overlap between reconstructed and ground-truth scenes. (iii) Semantic alignment. To assess alignment with input prompts, we render images from synthesized scenes and compute text-image similarity using CLIP [37] and BLIP [28], and image-image similarity using LPIPS (AlexNet) [67]. (iv) Aesthetic quality and functional plausibility. We evaluate aesthetic quality (AQ) and functional plausibility (FP) through human ratings from a user study with 16 participants and GPT-4o ratings. Methods are ranked based on averaged ordinal scores across a benchmark set of scenes, with lower ranks indicating better performance.
Table 1: Quantitative comparison of semantic alignment (CLIP, BLIP, LPIPS), aesthetic quality (AQ), and functional plausibility (FP).
Method CLIP ($\uparrow$) BLIP ($\uparrow$) LPIPS ($\downarrow$) AQ (4o/User) ($\downarrow$) FP (4o/User) ($\downarrow$)
@@
Table 2: Quantitative comparison of geometric fidelity and instance overlap on 3D-FRONT dataset [16, 17], and evaluation of the impact of the number of input view.
Method #Views Acc. ($\downarrow$) Comp. ($\downarrow$) CD ($\downarrow$) IoU ($\uparrow$)
ACDC 1 17.32 8.51 12.915 52.4
Ours 1 7.73 2.51 5.12 70.8
Ours 5 4.16 2.62 3.39 75.4
Ours 10 1.89 0.97 1.43 78.1
Ours 20 1.61 0.93 1.27 78.8
@@ 5.1 Comparison
Image-conditioned generation. Figure 4 shows a comparison with ACDC [9], a recent method for real-to-sim scene construction. Our system more effectively reconstructs spatial layouts and scene compositions, such as the tilted sofa in a living room, while better preserving the stylistic integrity of iconic works like Van Gogh’s Bedroom in Arles. In Table 1, our approach achieves the best image-image similarity score, demonstrating higher visual fidelity to the input images. For geometric accuracy and alignment at the instance level with ground truth, Table 2 shows that our approach achieves a lower Chamfer distance and higher IoU.
Image set as input. Unlike prior methods, which are typically restricted to single-view input, our framework naturally accommodates unstructured and unposed image collections. As illustrated in Figure 5, our system consolidates geometric cues from diverse viewpoints into a coherent 3D layout. This ability stems from the VLM’s integration with our spatial context, which provides a flexible representation for resolving spatial correspondences across views. We also evaluate the impact of the number of input views on performance. As shown in Table 2, increasing the number of views improves the precision of the reconstruction.